OpenStudy (chris215):

find f '(x) for f(x)=sqrt(sin(8x))

OpenStudy (anonymous):

I will use the chain rule to solve this problem: d/dx sqrt(sin(8x)) = (1/2)(sin(8x))^-(1/2) * d/dx(sin(8x)) = (1/2)(sin(8x))^-(1/2) * cos(8x) * d/dx 8x = (1/2)(sin(8x))^-(1/2) * cos(8x) * 8 = 4cos(8x)/sqrt(sin(8x)) Let me know if any steps are unclear, I am happy to go into more detail.

OpenStudy (chris215):

thank you!!